# Lie symmetry analysis of the two-Higgs-doublet model field equations

### v1

@@ 3 An illustrative example: real, scalar $\phi^{4}$ theory
The following Section may be skipped by readers familiar with Lie symmetry analysis of PDEs. The aim is here to illustrate the general theory of Sections 2.1-2.5 by a simple example, namely real, scalar $\phi^{4}$ theory.
@@ 3.1 $\phi^{4}$ theory in one dimension
where $\phi=\phi(x)$ is a scalar field depending on one free variable $x$, with a non-linear field equation
@@
The relevant prolongation (2.8) of the infinitesimal symmetry generator $X$ of (2.6) is
@@
where subscripts indicate partial derivatives. We then apply (3.1) on (3.2), and thereafter use the substitution
@@
everywhere we can, cf. the linearized symmetry condition (2.17), with the following determining equation as result,
@@
Here each coefficient of each distinct power of $\phi^{\prime}$ must vanish, for (3.1) to hold for any solution of the original field equation (3.2) (remember, $\xi$ and $\eta$ do not contain $\phi^{\prime}$), and we then get a larger system of determining equations,
@@
The determining equations (3.1) is a second order system of homogeneous, linear PDEs in the coefficients of the infinitesimal generator $X$, i.e. $\xi$ and $\eta$. It can be solved by elementary methods, and the solutions will depend on whether the parameters $m^{2}$ and $\lambda$ vanishes, since vanishing parameters will yield simpler equations with more solutions.
@@ 3.1.1 Massive $\phi^{4}$ theory
for a constant $c_{1}$, which give us the 1d Lie algebra defined by the generator
@@
which generates a translational symmetry $x\to x+\epsilon$, which is abstractly just a $\mathfrak{u}(1)$ symmetry. The generator (3.9) is here a strict variational symmetry, since
@@
cf. (2.21), since $\mathcal{L}$ doesn’t contain $x$ explicitly.
@@ 3.1.2 Massless $\phi^{4}$ theory
for arbitrary, real constants $c_{1}$ and $c_{2}$. The Lie symmetry (3.9) is hence enhanced to a 2d Lie algebra
@@
and we hence have the 2d, non-abelian Lie algebra $\mathfrak{a}(1)$. The generator $X_{1}$ will be strictly variational in the same manner as before. Concerning $X_{2}$,
@@
where $E(-3\mathcal{L})=-3E(\mathcal{L})$ is non-zero and of course proportional to the left hand side of the field equation. Hence the symmetry is non-variational, cf. (2.26), which means it is a symmetry of the field equation, but not the action. That $X_{2}$ actually is a symmetry of the field equation is confirmed by checking the linearized symmetry condition (2.17) for the field equation,
@@
which vanishes when $E(\mathcal{\mathcal{L}})=0$, i.e. when the field equation (3.2) (with $m=0$) holds, and $X_{2}$ is hence a symmetry of the equation. In 4d, the corresponding symmetry will be strictly variational, cf. (3.70).
@@ 3.1.3 Free, massive scalar theory
where the factors $(1/m)$ ensure we will get no occurences of the mass parameter $m$ in the commutator table. Here $X_{1}$ corresponds to translations and $X_{3}$ scaling symmetries. Some occurrences of the parameter $m^{-1}$ have here been absorbed in the constants $c_{j}$, which should be kept in mind if we want to explore the symmetries corresponding to the case $m=0$. The commutator table of (3.1.3) may then be calculated by applying SYM [45]
@@
We note that $\{X_{1},X_{7},X_{8}\}$, $\{X_{1},X_{5},X_{6}\}$ and $\{X_{1},X_{3},X_{7},X_{8}\}$ are subalgebras. The subalgebra $\{X_{3},X_{6}\}=\mathfrak{a}(1)$ is non-compact, so then the ambient algebra is non-compact as well. We will show that the Lie algebra generated by (3.1.3) is $\mathfrak{sl}(3)$ in Section (3.1.4). The only scalar symmetry $X_{3}$ has a prolongation which acts as
@@
and is hence not a variational symmetry, since the Euler operator of the result of (3.32) evidently is not zero: $E(2\mathcal{L})=2E(\mathcal{L})$ is a multiple of the Euler-Lagrange-expression, which is not identically zero. Of course, the expression vanishes when assuming the Euler-Lagrange equation hold, but we don’t assume this on the level of variational (also called ”off-shell”) symmetries.
@@ 3.1.4 Free, massless scalar theory
We again note that the algebra has an $\mathfrak{a}(1)=\{X_{1},X_{2}\}$ subalgebra, so the 8d algebra cannot be a compact algebra, since the subalgebra $\mathfrak{a}(1)$ is not compact. In fact, this algebra is $\mathfrak{sl}(3)$ (cf. Olver’s no 6.8 basis e.g. in [43]).
Lie himself showed that a second order equation with an 8d algebra, is always equivalent to the massless equation $\phi^{\prime\prime}=0$. The equation (3.2) with $\lambda=0,m^{2}\neq 0$ is equivalent to the massless equation $\hat{\phi}^{\prime\prime}=0$ through the point transformation (see e.g. [48])
@@
which makes the massless equation hold if and only if the massive equation holds, as long as $\cos(mx)\neq 0$. Note that the transformation (3.45) is not a symmetry, since the structure of the equation is not conserved.
Moreover, the algebra corresponding to (3.1.3) and (3.31) is $\mathfrak{sl}(3)$, just as for the case $m=0$. First, this is the only possible 8d algebra for any second order ODE. Second, the generators (3.1.3) may be transformed to generators yielding the exactly same commutator table as for $m=0$ (3.44), by a basis shift:
@@
we get a correspondence between the generators (3.1.3) related to $m\neq 0$, and the generators (3.1.4) corresponding to $m=0$. We will here denote the latter as hatted quantities, consistent with the transformation (3.45). For instance,
@@
where the last line refers to the generators of (3.1.4). Then the generators of (3.1.3) and (3.1.4) are connected through
@@
Then a calculation of the commutation table in the basis ${X}^{\prime}$, shows that the table now is identical to the one in (3.44), the $m=0$ case. Hence we have shown explicitly how the Lie symmetry algebra of the free, massive equation is the same as for the corresponding massless equation, namely $\mathfrak{sl}(3,\mathbb{R})$.
As in the massive case, the generator $X_{7}=\phi\partial_{\phi}$ will correspond to a non-variational symmetry, since its prolongation again will map $\mathcal{L}$ to a multiple of itself. The new scalar symmetry $X_{6}=\partial_{\phi}$ equals its own prolongation and will annihilate $\mathcal{L}=(1/2)(\phi^{\prime})^{2}$, and is thus a strict variational symmetry. The accordance with Theorem 1 is here trivial, since $V=0$.
@@ 3.2 $\phi^{4}$ theory in (3+1)d spacetime
Real, scalar $\phi^{4}$ theory in 4d has a Lagrangian density
@@
with a non-linear field equation
@@
We may now analyze the symmetries of (3.61), e.g. by applying MathLie [49] or SYM [45]. The single equation (3.61) here yields 19 determining equations.
@@ 3.2.1 The case $m,\lambda\neq 0$
In case $m,\lambda\neq 0$, the Lie symmetry algebra consists of 10 generators, the 6 first corresponding to the Lorentz algebra $\mathfrak{so}(1,3)$,
@@
where the three first generators are the boosts in $x_{1},x_{2}$ and $x_{3}$ directions, and the three last are the rotations about the $x_{1},x_{2}$ and $x_{3}$ axis, respectively. Together with the 4 translations
@@ 3.2.2 The case $m=0$ and $\lambda\neq 0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m=0$ and $\lambda\neq 0$ SYM [45] reveals the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced with the five generators [^8] Thanks to Stylianos Dimas for providing a tutorial Mathematica notebook demonstrating eqs. (3.2.1)–(3.2.2) and (3.2.3) using SYM.
@@
Scaling symmetries like $X_{11}$ are known for not being carried over to the quantum theory, due to the regulator introducing a scale. We will now demonstrate that the scaling symmetry generated by $X_{11}$ is, nevertheless, strictly variational, in contrast to the situation in the 1d case. The generator $X_{11}$ corresponds to infinitesimal symmetry transformations
@@
Then the derivative transforms as,
@@
to first order in the infinitesimal $\epsilon$, cf. (3.65). Then, suppressing $\epsilon$, $\delta\phi=-\phi$, $\delta x=x$, $\delta({\partial_{\mu}\phi})=-2{\partial_{\mu}\phi}$ and $\delta(d^{4}x)=4d^{4}x$, which gives a first order variation
@@
of the Lagrangian of $\phi^{4}$ theory. We get the same result by applying the 1-prolongation of $X_{11}$,
@@
to $\mathcal{L}$, since the definition (2.21) of a strict variational symmetry then holds,
@@
where $\xi^{\mu}=x^{\mu}$, in contrast to the 1d case (3.15). There are no scalar symmetries, just as the case was for the corresponding 1d theory of Section 3.1.2. This is in accordance with Corollary 1, which says the scalar, variational symmetries will be the same for the 1d and the 4d theories.
@@ 3.2.3 The case $m\neq 0$ and $\lambda=0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m\neq 0$ and $\lambda=0$ SYM [45] shows the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced with the generators
@@
Hence the symmetry algebra in this case becomes infinite dimensional, with $\mathfrak{iso}(1,3)$ as a finite subalgebra. The only scalar symmetry $X_{11}$ here is the same as for the 1d case in Section 3.1.3. By Corollary 1, the scalar, variational symmetries of the 4d theory must be the same as those of the 1d theory, which can easily be confirmed: As for the 1d case the prolongation of $X_{11}$ just maps $\mathcal{L}$ to a multiple of itself, and is hence again non-variational.
@@ 3.2.4 The case $m=\lambda=0$
In the case $m=\lambda=0$ the Poincaré algebra $\mathfrak{iso}(1,3)$ is enhanced with both sets of generators given by (3.2.2) and (3.2.3), and hence the algebra in this case again becomes infinite dimensional. The symmetry $X_{11}=\phi\partial_{\phi}$ will still map $\mathcal{L}$ to a multiple of itself, as in Section 3.1.4, and is hence non-variational as before. The 4d, scalar symmetry is hence of the same nature as in the 1d case, in harmony with Corollary 1.
No divergence symmetries were found in the analysis in Section 3, but they occur e.g. in scalar gauge singlet extensions of the SM [50]. We finally also note all three cases $m=0$, $\lambda=0$ and $\lambda=m=0$ may be imposed by symmetries, since all cases are corresponding to distinct symmetries.
@@
where covariant derivatives and the gauge boson field strength tensors are given by (2.52)-(2.54). The most general, renormalizable 2HDM potential may be written as
@@
where ”h.c.” means Hermitian conjugate, i.e. conjugate transpose, with complex parameters $m_{12}^{2}$, $\lambda_{5}$, $\lambda_{6}$ and $\lambda_{7}$, other parameters real. By a basis shift (i.e. a reparametrization) of the doublets,
@@
with [^9] Or with $U\in\mathsf{SU}(2)$, if we absorb the complex phase of $\mathsf{U}(2)$ into $\mathsf{U}(1)_{Y}$ hypercharge symmetry. $U\in\mathsf{U}(2)$, we may diagonalize the mass-squared matrix and remove the complex phase of e.g. $\lambda_{5}$, which means
@@
The 24 Euler-Lagrange equations $E(\mathcal{L})=0$ of (4.1) are of the form
@@
These equations will be variants of interacting (i.e. nonlinear) Klein-Gordon and Proca equations.

### v2

@@ 3 An illustrative example: real, scalar $\phi^{4}$ theory
The following section may be skipped by readers familiar with Lie symmetry analysis of PDEs. The aim is here to illustrate the general theory of Sections 2.1-2.5, using a simple example, namely real, scalar $\phi^{4}$ theory.
@@ 3.1 $\phi^{4}$ theory in one dimension
where $\phi=\phi(x)$ is a scalar field that depends on one free variable $x$, with a nonlinear field equation
@@
In the standard real $\phi^{4}$ model (3.1) one usually imposes a discrete $\mathbb{Z}_{2}$ symmetry $\phi\to-\phi$ in order to forbid linear and cubic terms. This yields a simple, renormalizable model with a symmetric vacuum structure and a potential containing only the even powers $\phi^{2}$ and $\phi^{4}$. The relevant prolongation (2.10) of the infinitesimal symmetry generator $X$ in (2.8) is
@@
where subscripts indicate partial derivatives. We then apply (3.1) to (3.2), and thereafter use the substitution
@@
everywhere we can, cf. the linearized symmetry condition (2.19), with the following determining equation as result:
@@
Here, each coefficient of each distinct power of $\phi^{\prime}$ must vanish for (3.1) to hold for any solution of the original field equation (3.2) (remember, $\xi$ and $\eta$ do not contain $\phi^{\prime}$), and we then obtain a larger system of determining equations:
@@
The determining equations (3.1) are a second-order system of homogeneous, linear PDEs in the coefficients of the infinitesimal generator $X$, that is, $\xi$ and $\eta$. The system can be solved by elementary methods, and the solutions will depend on whether the parameters $m^{2}$ and $\lambda$ vanish, as the vanishing parameters will yield simpler equations with more solutions.
@@ 3.1.1 Massive $\phi^{4}$ theory
for a constant $c_{1}$, which gives us the 1d Lie algebra defined by the generator
@@
which generates the translational symmetry $x\to x+\epsilon$ and abstractly spans the one-dimensional Lie algebra $\mathbb{R}$. The generator (3.9) is here a strict variational symmetry, because
@@
cf. (2.23), since $\mathcal{L}$ does not contain $x$ explicitly.
@@ 3.1.2 Massless $\phi^{4}$ theory
for arbitrary, real constants $c_{1}$ and $c_{2}$. Hence, the Lie symmetry (3.9) is enhanced to a 2d Lie algebra
@@
and we therefore have the 2d, non-abelian Lie algebra $\mathfrak{a}(1)$. The generator $X_{1}$ is strictly variational in the same manner as before. Concerning $X_{2}$,
@@
where $E(-3\mathcal{L})=-3E(\mathcal{L})$ is non-zero and is proportional to the left-hand side of the field equation. Hence, the symmetry is non-variational, cf. (2.28), which means that it is a symmetry of the field equation, but not the action. That $X_{2}$ is actually a symmetry of the field equation is confirmed by checking the linearized symmetry condition (2.19) for the field equation:
@@
which vanishes when $E(\mathcal{\mathcal{L}})=0$, that is, when the field equation (3.2) (with $m=0$) holds, and $X_{2}$ is hence a symmetry of the equation. In 4d, the corresponding symmetry will be strictly variational, cf. (3.70).
@@ 3.1.3 Free, massive scalar theory
where the factors $(1/m)$ ensure that there are no occurrences of the mass parameter $m$ in the commutator table. Here, $X_{1}$ corresponds to translations and $X_{3}$ corresponds to scaling symmetries. Some occurrences of the parameter $m^{-1}$ have been absorbed in the constants $c_{j}$, which should be kept in mind if we want to explore the symmetries corresponding to the case $m=0$. The commutator table of (3.1.3) can then be calculated by applying SYM [43]
@@
Note that $\{X_{1},X_{7},X_{8}\}$, $\{X_{1},X_{5},X_{6}\}$ and $\{X_{1},X_{3},X_{7},X_{8}\}$ are subalgebras. The subalgebra $\{X_{3},X_{6}\}=\mathfrak{a}(1)$ is non-compact; thus, the ambient algebra is non-compact as well. We show that the Lie algebra generated by (3.1.3) is $\mathfrak{sl}(3)$ in Section (3.1.4). The only scalar symmetry $X_{3}$ has a prolongation that acts as
@@
and is hence not a variational symmetry, because the Euler operator of the result of (3.32) is evidently not zero: $E(2\mathcal{L})=2E(\mathcal{L})$ is a multiple of the Euler-Lagrange-expression, which is not identically zero. Of course, the expression vanishes when assuming that the Euler-Lagrange equation holds; however we do not assume this on the level of variational (also called ”off-shell”) symmetries.
@@ 3.1.4 Free, massless scalar theory
We again note that the algebra has an $\mathfrak{a}(1)=\{X_{1},X_{2}\}$ subalgebra, so the 8d algebra cannot be a compact algebra, because the subalgebra $\mathfrak{a}(1)$ is not compact. This algebra is $\mathfrak{sl}(3)$ (cf. Olver’s no 6.8 basis, e.g. in [46]).
Lie himself showed that a second-order equation with an 8d algebra, is always equivalent to the massless equation $\phi^{\prime\prime}=0$. Equation (3.2) with $\lambda=0,m^{2}\neq 0$ is equivalent to the massless equation $\hat{\phi}^{\prime\prime}=0$ through the point transformation (see e.g. [47])
@@
which makes the massless equation hold if and only if the massive equation holds, as long as $\cos(mx)\neq 0$. Note that the transformation (3.45) is not a symmetry, because the structure of the equation is not conserved.
Moreover, the algebra corresponding to (3.1.3) and (3.31) is $\mathfrak{sl}(3)$, which is the same algebra as in the case $m=0$. First, it is the only possible 8d algebra for any second-order ODE. Second, the generators (3.1.3) may be transformed to generators yielding the same commutator table as for $m=0$ (3.44), by a basis shift:
@@
we obtain a correspondence between the generators (3.1.3) related to $m\neq 0$, and the generators (3.1.4) corresponding to $m=0$. Here, we will denote the latter as hatted quantities, consistent with the transformation (3.45). For instance,
@@
where the last line refers to the generators of (3.1.4). Subsequently, the generators of (3.1.3) and (3.1.4) are connected through
@@
Then, a calculation of the commutation table in the basis ${X}^{\prime}$, shows that the table is identical to that in (3.44), the $m=0$ case. Hence, we have shown explicitly how the Lie symmetry algebra of the free, massive equation is the same as that of the corresponding massless equation, namely $\mathfrak{sl}(3,\mathbb{R})$.
As in the massive case, the generator $X_{7}=\phi\partial_{\phi}$ will correspond to a non-variational symmetry because its prolongation will again map $\mathcal{L}$ to a multiple of itself. The new scalar symmetry $X_{6}=\partial_{\phi}$ equals its own prolongation and annihilates $\mathcal{L}=(1/2)(\phi^{\prime})^{2}$, and is thus a strict variational symmetry. The accordance with Theorem 1 is trivial, because $V=0$.
@@ 3.2 $\phi^{4}$ theory in (3+1)d spacetime
Real scalar $\phi^{4}$ theory in 4d has a Lagrangian density
@@
with a nonlinear field equation
@@
for $\phi=\phi(x)$. We may now analyze the symmetries of (3.61), for example by applying MathLie [48] or SYM [43]. Equation (3.61) then yields 19 determining equations.
@@ 3.2.1 The case $m,\lambda\neq 0$
In the case $m,\lambda\neq 0$, the Lie symmetry algebra consists of 10 generators, the first six of which correspond to the Lorentz algebra $\mathfrak{so}(1,3)$,
@@
where the first three generators are the boosts in $x_{1},x_{2}$ and $x_{3}$ directions, and the last three are the rotations about the $x_{1},x_{2}$ and $x_{3}$ axes, respectively. Together with the four translations
@@ 3.2.2 The case $m=0$ and $\lambda\neq 0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m=0$ and $\lambda\neq 0$ SYM [43] reveals that the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced with the five generators [^8] Thanks to Stylianos Dimas for providing a tutorial Mathematica notebook demonstrating equations (3.2.1)–(3.2.2) and (3.2.3)–(3.72) using SYM.
@@
Scaling symmetries such as $X_{11}$ are known not to be preserved in the quantum theory, due to the regulator introducing a scale. We now demonstrate that the scaling symmetry generated by $X_{11}$ is, nevertheless, strictly variational, in contrast to the situation in the 1d case. The generator $X_{11}$ corresponds to infinitesimal symmetry transformations
@@
Then the derivative transforms as
@@
to first order in the infinitesimal parameter $\epsilon$, cf. (3.65). Then, suppressing $\epsilon$, $\delta\phi=-\phi$, $\delta x=x$, $\delta({\partial_{\mu}\phi})=-2{\partial_{\mu}\phi}$ and $\delta(d^{4}x)=4d^{4}x$, which gives a first-order variation
@@
of the Lagrangian of $\phi^{4}$ theory. We obtain the same result by applying the 1-prolongation of $X_{11}$,
@@
to $\mathcal{L}$, because the definition (2.23) of a strict variational symmetry then holds,
@@
where $\xi^{\mu}=x^{\mu}$, in contrast to the 1d case (3.15). There are no scalar symmetries, as in the case of the corresponding 1d theory in Section 3.1.2. This is in accordance with Corollary 1, which states that scalar variational symmetries will be the same for the 1d and 4d theories.
@@ 3.2.3 The case $m\neq 0$ and $\lambda=0$
In the case of $\phi^{4}$ theory in (3+1)d spacetime with $m\neq 0$ and $\lambda=0$ SYM [43] shows the symmetry algebra $\mathfrak{iso}(1,3)$ is enhanced by the generators
@@
Hence the symmetry algebra in this case becomes infinite dimensional, with $\mathfrak{iso}(1,3)$ as a finite subalgebra. The only scalar symmetry $X_{11}$ here is the same as for the 1d case in Section 3.1.3. By Corollary 1, the scalar, variational symmetries of the 4d theory must be the same as those of the 1d theory, which can easily be confirmed: As for the 1d case, the prolongation of $X_{11}$ maps $\mathcal{L}$ to a multiple of itself, and is hence again non-variational.
@@ 3.2.4 The case $m=\lambda=0$
In the case $m=\lambda=0$, the Poincaré algebra $\mathfrak{iso}(1,3)$ is enhanced by both sets of generators given in (3.2.2) and (3.2.3); hence the algebra once again becomes infinite-dimensional. The symmetry $X_{11}=\phi\partial_{\phi}$ will still map $\mathcal{L}$ to a multiple of itself, as in Section 3.1.4, and is hence non-variational as before. The 4d, scalar symmetry is thus of the same nature as in the 1d case, in harmony with Corollary 1.
No divergence symmetries were found in the analysis in Section 3, but they occur e.g. in scalar gauge singlet extensions of the SM [49]. Finally, we also note that all three cases, $m=0$, $\lambda=0$ and $\lambda=m=0$ may be imposed by symmetries, as all cases correspond to distinct symmetries.
@@
where the covariant derivatives and gauge boson field strength tensors are given by (2.55)-(2.57). The most general, renormalizable 2HDM potential can be written as
@@
where ”h.c.” means Hermitian conjugate, i.e. conjugate transpose, with complex parameters $m_{12}^{2}$, $\lambda_{5}$, $\lambda_{6}$ and $\lambda_{7}$, other parameters real. Applying a basis shift (i.e. a reparametrization) of the doublets,
@@
with [^9] Or with $U\in\mathsf{SU}(2)$, if we absorb the complex phase of $\mathsf{U}(2)$ into $\mathsf{U}(1)_{Y}$ hypercharge symmetry. $U\in\mathsf{U}(2)$, we can diagonalize the mass-squared matrix and remove the complex phase of e.g. $\lambda_{5}$, which means that
@@
The 24 Euler-Lagrange equations $E(\mathcal{L})=0$ in (4.1) are of the form
@@
These equations will be variants of the interacting (i.e., nonlinear) Klein-Gordon and Proca equations.